OpenStudy (anonymous):

Otto mixed spices to create his own Italian seasoning. He used twenty percent garlic powder. Which of the following shows the percent equation and solution for how many tablespoons of garlic powder he uses to create sixteen tablespoons of the seasoning? a equals zero point two times sixteen; a equals three point two sixteen equals zero point two times a; a equals eighty a equals zero point two times sixteen; a equals eighty sixteen equals zero point two times a; a equals three point two

OpenStudy (triciaal):

per cent means per 100 20/100 of what amount,a = 16 solve for a

OpenStudy (triciaal):

same as 20/100 = 16/a